. if Qd = a/b - cP and Qs = d/e + fP, in market equilibrium (Qd = Qs) what is P, Qd and Qs?



1
Expert's answer
2021-02-01T09:01:44-0500

Equating "Q_d" and "Q_s", we get:


"\dfrac{a}{b}-cP=\dfrac{d}{e}+fP,""P=\dfrac{ae-bd}{be(c+f)}."

Substitunig "P" into the "Q_d" and "Q_s", we get:


"Q_d=\dfrac{a}{b}-c\dfrac{(ae-bd)}{be(c+f)},""Q_s=\dfrac{d}{e}+f\dfrac{(ae-bd)}{be(c+f)}."